Assignment I: Assume you have $50,000 to invest over the next five years. At the beginning of each year you can invest the available money in one, two or three-year time deposits. The bank pays 1.1% interest on one-year deposits, 2.1% interest (total) on two-year time deposits, and 3.2% interest (total) on three-year deposits. For instance, if you invest $1000 at the beginning of year 1 in one-year deposits, your return at the end of year 1, or beginning of year 2, would be 1.01 1000-$1010. This amount ($1010) wll be invested at the beginning of year 2. If you invest the money that is available at the beginning of every year, formulate a linear programming to maximize your total cash on hand at the end of the fifth year A: Assume that you can invest at the beginning of each year any value in the available your linear programming model. (6 marks) B: Solve your investment LP model using Excel-Solver and obtain the optimal solution.
Answer #1

A)

A11 Amount invested for 1 year in 1 year

A12 Amount invested for 2 year in 1 year

A13 Amount invested for 3 year in 1 year

A21 Amount invested for 1 year in 2 year

A22 Amount invested for 2 year in 2 year

A23 Amount invested for 3 year in 2 year

A31 Amount invested for 1 year in 3 year

A32 Amount invested for 2 year in 3 year

A33 Amount invested for 3 year in 3 year

A41 Amount invested for 1 year in 4 year

A42 Amount invested for 2 year in 4 year

A51 Amount invested for 1 year in 5 year

Total amount, T = A33*(1+1.1%)^3+A42*(1+2.1%)^2+A51*(1+3.1%)

Objective is Max T subject to

Constraints

A11 + A12 + A13 <= 50000

A21+A22+A23<= A11*1.011

A31+A32+A33<= A21*1.011+A12*1.021^2

A41+A42<= A13*1.031^3+A22*1.021^2+A31*1.011

A51<= A41*1.011

A11 A12 A13 Amount invested for Amount invested for Amount invested for Amount invested for Amount invested for Amount invested for Amount invested for Amount invested for Amount invested for Amount invested for Amount invested for Amount invested for r in year in year in r in r in year in year in year in year in year in year in year in year S0000.00 year year year year year A31 A32 A33 A41 year 54795.64 A51 year Total amount, T- A33(1+1.1%)^3+A42(1+2.1%)2+A51*(1+3.1%) 57121.22129 Objective is Max T subject to Constraints A11 +A12 +A13 A21+A22+A23 A31+A32+A33 A41+A42 AS1 50000.00 0.00 0.00 54795.64 0.00 50000.00 0.00 0.00 54795.64 0.00 50000 A11 1.011 A21 1.011+A121.0212 A13 1.0313+A22 1.021 2+A31*1.031 A41*1.011